# Your Title

### v1

@@ 3.3 Adaptive Motion-Driven Mechanism
Equipped with the selected keypoints $\mathcal{K}_{t}$ at current timestep, the next step is to determine which neighboring points are controlled by these keypoints, and apply their transformations to drive the motion of the controlled neighboring points. Previous works [43, 44] employ k-nearest neighbor (KNN) [45] search to predict the motion of each Gaussian points, showing advanced results in monocular synthetic video reconstruction, but they do not fully consider unnecessary modeling in static region and motion scale difference, which leads to computational redundancy and inaccurate representation.
In contrast, we propose an adaptive motion-driven mechanism that enables each keypoint to drive neighboring points through a spatial influence field, as illustrated Fig. 2 (c). Specifically, motivated by [9], for each keypoints $\mathcal{K}_{t}^{i}$ at $t$ timestep, we initialize a quaternion $q_{adap}^{i}\in\mathbb{R}^{4}$ and a scaling vector $s_{adap}^{i}\in\mathbb{R}^{3}$ to compute the spatial influence field $\Sigma_{adap}^{i}\in\mathbb{R}^{3\times 3}$. For a neighboring Gaussian point $G_{j}$ with position $\mu_{j}$, its distance to keypoint $\mathcal{K}_{t}^{i}$ is given by $\mu_{ij}=\mu_{j}-\mu_{\mathcal{K}_{t}^{i}}$. The influence weight is then computed as:
$w_{ij}=\exp\left(-\frac{1}{2}\mu_{ij}^{\top}(\Sigma_{adap}^{i})^{-1}\mu_{ij}\right)$ (6)
If $w_{ij}$ exceeds a predefined threshold $\tau_{adap}$, the Gaussian $G_{j}$ is considered to be controlled by keypoint $\mathcal{K}_{t}^{i}$.
$\text{If}\quad w_{ij}\geq\tau_{adap},\quad G_{j}\in\mathcal{C}_{t}^{i}$ (7)
where $\mathcal{C}_{t}^{i}$ denotes the set of Gaussians controlled by keypoint $\mathcal{K}_{t}^{i}$.
To model motion, each keypoint $\mathcal{K}_{t}^{i}$ is further assigned a learnable translation offset $\Delta\mu_{\mathcal{K}}^{i}\in\mathbb{R}^{3}$ and a rotation represented by a quaternion $\Delta q_{\mathcal{K}}^{i}\in\mathbb{R}^{4}$. For a Gaussian $G_{j}$ controlled by multiple keypoints $\{\mathcal{K}_{t}^{i}\}_{i\in\mathcal{I}_{j}}$, its overall motion is computed by aggregating the motions of its associated keypoints, weighted by their influence scores ${w}_{ij}$:
$\Delta\mu_{j}^{t}=\sum_{i\in\mathcal{I}_{j}}{w}_{ij}\cdot\Delta\mu_{\mathcal{K}}^{i},\quad\Delta q_{j}^{t}=\sum_{i\in\mathcal{I}_{j}}{w}_{ij}\cdot\Delta q_{\mathcal{K}}^{i}$ (8)
where $\Delta\mu_{j}^{t}$ and $\Delta q_{j}^{t}$ indicate the motion of Gaussian $j$ at $t$ timestep.
@@ 3.4 Error-Aware Corrector
By using keypoints to drive scene motion, we model the transformation of Gaussian points from the previous frame to the current frame with an extremely compact parameters. Nevertheless, keypoints-based motion controlling only supports to represent rigid motion effectively and faces challenge to handle non-rigid motion and novel objects emergence, which results in error accumulation across frames.
A straightforward solution to mitigate error accumulation and ensure accurate long-term FVV representations is to separate the video into frame groups and update the attributes of all Gaussians at key frames. However, this strategy would lead to a substantial of unnecessary parameters updating, since most of parameters are already correctly representing the scene and do not require modification. To mitigate error accumulation in a compact and efficient manner, we propose an error-aware corrector strategy that only finetunes the Gaussians with detected errors, significantly decreasing storage demands and promoting more accurate scene reconstruction, as illustrated in Fig. 2 (d).
Specifically, given a video sequence, we select a key frame every $s$ frames, forming the key frame sequence $\{{f}_{s},{f}_{2s},\dots,{f}_{ns}\}$, as shown in Fig. 2 (a). The remaining frames are reconstructed by keypoints driven. During key frame reconstruction, given the attributes of a Gaussian point at previous timestep $\theta_{i}^{t-1}:(\mu_{i}^{t-1},q_{i}^{t-1},s_{i}^{t-1},\sigma_{i}^{t-1},sh_{i}^{t-1})$, we introduce a set of learnable parameters $\Delta\theta_{i}^{t}$ to model the attribute residuals. To identify which Gaussian points require correction, we predict a learnable mask $m_{i}$ for each point. A sigmoid function is used to map $m_{i}$ to the range $(0,1)$, which refers as a soft mask:
@@
Similar to [21, 22], the soft mask is binarized into a hard mask using a predefined threshold $\phi_{{thres}}$, where the non-differentiable binarization is handled with the straight-through estimator (STE) to enable gradient flow, represented as:
@@
$\theta_{i}^{t}=\theta_{i}^{t-1}\oplus m_{i}^{hard}*\Delta\theta_{i}^{t}$ (11)
where $\oplus$ denotes the attribute-specific update operation. Meanwhile, we define a optimized function to regulate the perceptual error while encouraging sparse residual updates:
@@
After optimization for the current key frame, only the attribute residuals $\hat{\Delta\theta^{t}}=\{\Delta\theta^{t}|\mathcal{M}^{hard}=1\}$ and the hard mask set $\mathcal{M}^{{hard}}$ need to be stored and transmitted, minimizing the required data redundancy and transmission overhead.
@@ 3.5 Optimization and Compression
For the first frame optimization, we employ COLMAP [27] to generate the initial point cloud and follow the pipeline of 3DGStream [13]. The optimization for both the first frame and non-key frames is supervised by the reconstruction loss $\mathcal{L}_{{recon}}$, which is composed by an $L_{1}$-norm loss $\mathcal{L}_{1}$ and a D-SSIM loss $\mathcal{L}_{D-SSIM}$ [24]:
@@
Table 1: Quantitative comparisons on Neural 3D Video (N3DV) [14] and MeetRoom datasets [25].
Dataset Category Method PSNR (dB) $\uparrow$ Storage (MB) $\downarrow$ Training (sec) $\downarrow$ Rendering (FPS) $\uparrow$
N3DV Offline NeRFPlayer [29] 30.69 17.10 72 0.05
HyperReel [30] 31.10 1.20 104 2.00
4DGS [11] 31.15 0.13 8 34
SpaceTime [12] 32.05 0.67 20 140
Online StreamRF [25] 30.68 31.4 15 8.3
TeTriRF [36] 30.43 0.06 39 4
3DGStream [13] 31.67 7.80 8.5 261
QUEEN-s [32] 31.89 0.68 4.65 345
QUEEN-l [32] 32.19 0.75 7.9 248
ComGS-s (ours) 31.87 0.049 37 91
ComGS-l (ours) 32.12 0.106 43 147
MeetRoom Static I-NGP [35] 28.10 48.2 66 4.1
3DG-S [9] 31.31 21.1 156 571
Online StreamRF [25] 26.72 9.0 10.2 10
3DGStream [13] 30.79 4.1 4.9 350
ComGS-s (ours) 31.49 0.028 28.3 98
Figure 3: Quantitative comparison. We visualize our method and other online FVV methods on N3DV [14] and MeetRoom [25] dataset.
Table 2: Ablation study on proposed components. Flame Steak and Flame Salmon are from the N3DV dataset.
@@
PSNR (dB)$\uparrow$ Storage (KB)$\downarrow$ PSNR (dB)$\uparrow$ Storage (KB)$\downarrow$
@@ 4.1 Experimental Setup
We evaluate our method on two widely-used public benchmark datasets. (1) Neural 3D Video (N3DV) [14] consists of six indoor video sequences captured by 18 to 21 viewpoints. (2) Meet Room [25] comprises four indoor scenes recorded with a 13 cameras multi-view system. In both of two datasets, we employ the first view for testing. Our method is implemented on an NVIDIA A100 GPU. We train 150 epochs for non-key frames reconstruction and 1000 epochs for key frames fine-tuning. We measure the visual quality of rendered images by average PSNR, required storage, rendering FPS and training time. More implement details are provided in the Appendix.
@@ 4.2 Quantitative Comparisons
We conduct quantitative comparisons on existing online methods including StreamRF [25], TeTriRF [36], 3DGStream [13] and QUEEN [32], as well as the SOTA offline FVV approaches [12, 11, 29, 30] on N3DV and Meetroom (Tab. 1). Our method is evaluated in two variants: ComGS-s (small) and ComGS-l (large), using key frame intervals of $s{=}10$ and $s{=}2$, respectively.
Tab. 1 shows that our ComGS achieves competitive results among existing online FVV methods on N3DV dataset. Notably, ComGS-s achieves a substantial reduction in storage by 159×159$\times$159 × compared to 3DGStream and 14×14$\times$14 × compared to QUEEN. This advantage enables real-time transmission in limited bandwidth and enhances the overall user viewing experience. On MeetRoom dataset, our method outperforms 3DGStream, obtaining +0.7dB PSNR and 146 $\times$ smaller size. Our advantages are mainly due to two factor: 1) using keypoint as a shared representation requires transmitting only a small number of keypoint attributes; and 2) the error-aware correction module effectively rectifies regions with scene inaccuracies using minimal additional parameters. In the Appendix, the quantitative results are provided for each scene to offer a more detailed comparison.
@@ 4.4 Ablation Study
To validate the effectiveness of our proposed methods, we ablate three components of ComGS framework in Tab. 2. In the Experiment 1, we adopt a random chosen techniques for keypoint selection. However, since the selection is not guided by region regions, it may result in ineffective modeling of static areas and inadequate representation of scenes, which leads to a slight degradation in PSNR. Experiment 2 model scene motion only using the selected keypoints, without applying the adaptive motion-driven mechanism for neighboring points. The resulting drop in reconstruction quality demonstrates the importance of collaborative modeling in keypoint-driven regions for accurate motion representation. In the Experiment 3, we reconstruct the FVV only relying on the key frames correction, which indicates that keypoint-driven motion representation is crucial for our method. Experiment 4 disable the error-aware correction in key frames reconstruction. This leads to a significant performance drop, demonstrating that error-aware correction in key frames would solve the error accumulation across frames.
@@
To further investigate the role of keypoint-driven motion representation, we visualize the selection and driven process in Fig. 4. The top row shows that keypoints are predominantly selected in motion regions, such as the human body and moving objects. The bottom row highlights the adaptively controlled areas for neighboring points, which similarly focus on regions with significant motion (e.g., the person and the dog). Fig. 5 visualizes Gaussians updated region using farthest keypoint selection [44], random keypoint selection and our method, respectively, which demonstrates that our method accurately captures motion-intensive areas. These results indicate that ComGS can effectively leverage the locality and consistency of scene motion.
We also evaluate a KNN-based method [45] for selecting neighboring points around keypoints (Tab. LABEL:Tab:_ablation_knn). This approach shows inferior performance, as it does not distinguish between static and motion regions, leading to redundant modeling and poor adaptation to varying motion scales.

### v2

@@ 3.3 Adaptive Motion-Driven Mechanism
Equipped with the selected keypoints $\mathcal{K}_{t}$ at current timestep, the next step is to determine which neighboring points are controlled by these keypoints, and apply their transformations to drive the motion of the controlled neighboring points. Previous works [huang2024sc, wan2024superpoint] employ k-nearest neighbor (KNN) [peterson2009k] search to predict the motion of each Gaussian points, showing advanced results in monocular synthetic video reconstruction, but they do not fully consider unnecessary modeling in static region and motion scale difference, which leads to computational redundancy and inaccurate representation.
In contrast, we propose an adaptive motion-driven mechanism that enables each keypoint to drive neighboring points through a spatial influence field, as illustrated Fig. 2 (c). Specifically, motivated by [kerbl20233d], for each keypoints $\mathcal{K}_{t}^{i}$ at $t$ timestep, we initialize a quaternion $q_{adap}^{i}\in\mathbb{R}^{4}$ and a scaling vector $s_{adap}^{i}\in\mathbb{R}^{3}$ to compute the spatial influence field $\Sigma_{adap}^{i}\in\mathbb{R}^{3\times 3}$. For a neighboring Gaussian point $G_{j}$ with position $\mu_{j}$, its distance to keypoint $\mathcal{K}_{t}^{i}$ is given by $d_{ij}=\mu_{j}-\mu_{\mathcal{K}_{t}^{i}}$. The influence weight is then computed as:
$w_{ij}=\exp\left(-\frac{1}{2}d_{ij}^{\top}(\Sigma_{adap}^{i})^{-1}d_{ij}\right)$ (7)
If $w_{ij}$ exceeds a predefined threshold $\tau_{adap}$, the Gaussian $G_{j}$ is considered to be controlled by keypoint $\mathcal{K}_{t}^{i}$:
$\mathcal{C}_{t}^{i}=\{G_{j}\ |\ w_{ij}\geq\tau_{adap}\}$ (8)
where $\mathcal{C}_{t}^{i}$ denotes the set of Gaussian points controlled by keypoint $\mathcal{K}_{t}^{i}$.
To model motion, each keypoint $\mathcal{K}_{t}^{i}$ is further assigned a learnable translation offset $\Delta\mu_{\mathcal{K}^{i}_{t}}\in\mathbb{R}^{3}$ and a rotation represented by a quaternion $\Delta q_{\mathcal{K}^{i}_{t}}\in\mathbb{R}^{4}$. For a Gaussian $G_{j}$ controlled by multiple keypoints $\{\mathcal{K}_{t}^{i}\}_{i\in\mathcal{I}_{t}^{j}}$, its overall motion is computed by aggregating the motions of its associated keypoints, weighted by their influence scores ${w}_{ij}$:
$\Delta\mu^{j}_{t}=\sum_{i\in\mathcal{I}_{t}^{j}}{w}_{ij}\cdot\Delta\mu_{\mathcal{K}^{i}_{t}},\quad\Delta q^{j}_{t}=\sum_{i\in\mathcal{I}_{t}^{j}}{w}_{ij}\cdot\Delta q_{\mathcal{K}^{i}_{t}}$ (9)
where $\Delta\mu^{j}_{t}$ and $\Delta q^{j}_{t}$ indicate the transformation of Gaussian $j$ at $t$ timestep, and $\mathcal{I}_{t}^{j}$ represents the set of keypoints that control the motion of Gaussian $j$.
@@ 3.4 Error-Aware Corrector
By using keypoints to drive scene motion, we model the transformation of Gaussian points from the previous frame to the current frame with an extremely compact parameters. Nevertheless, keypoint-based motion controlling only supports to represent rigid motion effectively and faces challenge to handle non-rigid motion and novel objects emergence, which results in error accumulation across frames.
A straightforward solution to mitigate error accumulation and ensure accurate long-term FVV representations is to separate the video into frame groups and update the attributes of all Gaussian points at key frames. However, this strategy would lead to a substantial of unnecessary parameters updating, since most of parameters are already correctly representing the scene and do not require modification. To mitigate error accumulation in a compact and efficient manner, we propose an error-aware corrector strategy that only finetunes the Gaussian points with detected errors, significantly decreasing storage demands and promoting more accurate scene reconstruction, as illustrated in Fig. 2 (d).
Specifically, given a video sequence, we select a key frame every $s$ frames, forming the key frame sequence $\{{f}_{s},{f}_{2s},\dots,{f}_{ns}\}$, as shown in Fig. 2 (a). Note that in this paper, key frames are used for error correction, and only the first frame of the video sequence is independently reconstructed. The remaining frames are reconstructed by keypoints driven. During key frame reconstruction, given the attributes of a Gaussian point at previous timestep $\theta_{i}^{t-1}:(\mu_{i}^{t-1},q_{i}^{t-1},s_{i}^{t-1},\sigma_{i}^{t-1},c_{i}^{t-1})$, we introduce a set of learnable parameters $\Delta\theta_{i}^{t}$ to model the attribute residuals. To identify which Gaussian points require correction, we predict a learnable mask $m_{i}$ for each point. A sigmoid function is used to map $m_{i}$ to the range $(0,1)$, which refers as a soft mask:
@@
Similar to [lee2024compact, wang2025end], the soft mask is binarized into a hard mask using a predefined threshold $\phi_{{thres}}$, where the non-differentiable binarization is handled with the straight-through estimator (STE) to enable gradient flow, represented as:
@@
$\theta_{i}^{t}=\theta_{i}^{t-1}+m_{i}^{hard}\Delta\theta_{i}^{t}$ (12)
Meanwhile, we define a optimized function to regulate the perceptual error while encouraging sparse residual updates:
@@
where $N$ is the number of all Gaussian points. After optimization for the current key frame, only the attribute residuals $\hat{\Delta\theta^{t}}=\{\Delta\theta^{t}_{i}|m^{hard}_{i}=1\}$ and the hard mask set $\mathcal{M}^{{hard}}=\{m^{hard}_{i}|i=1,2,...,N\}$ need to be stored and transmitted, minimizing the required data redundancy and transmission overhead.
@@ 3.5 Optimization and Compression
For the first frame optimization, we employ COLMAP [schonberger2016structure] to generate the initial point cloud and follow the pipeline of 3DGStream [sun20243dgstream]. The optimization for both the first frame and non-key frames is supervised by the reconstruction loss $\mathcal{L}_{{recon}}$, which is composed by an $L_{1}$-norm loss $\mathcal{L}_{1}$ and a D-SSIM loss $\mathcal{L}_{D-SSIM}$ [wang2004image]:
@@
Table 1: Quantitative comparisons on Neural 3D Video (N3DV) [li2022neural] and MeetRoom datasets [li2022streaming].
Dataset Category Method PSNR SSIM $\uparrow$ LPIPS $\downarrow$ Storage Training Rendering
(dB) $\uparrow$ (MB) $\downarrow$ (sec) $\downarrow$ (FPS) $\uparrow$
N3DV Offline NeRFPlayer [song2023nerfplayer] 30.69 0.932 0.209 17.10 72 0.05
HyperReel [attal2023hyperreel] 31.10 0.928 - 1.20 104 2.00
4D-GS [wu20244d] 31.15 0.964 0.149 0.13 8 34
SpaceTime [li2024spacetime] 32.05 0.948 - 0.67 20 140
Online StreamRF [li2022streaming] 30.68 - - 31.4 15 8.3
TeTriRF [wu2024tetrirf] 30.43 0.906 0.248 0.06 39 4
3DGStream [sun20243dgstream] 31.67 0.941 0.140 7.80 8.5 261
QUEEN-s [girish2025queen] 31.89 0.945 0.139 0.68 4.65 345
QUEEN-l [girish2025queen] 32.19 0.946 0.136 0.75 7.9 248
ComGS-s (ours) 31.87 0.943 0.132 0.049 37 91
ComGS-l (ours) 32.12 0.945 0.129 0.106 43 147
MeetRoom Static I-NGP [jiang2023instant] 28.10 - - 48.2 66 4.1
3DG-S [kerbl20233d] 31.31 - - 21.1 156 571
Online StreamRF [li2022streaming] 26.72 - - 9.0 10.2 10
3DGStream [sun20243dgstream] 30.79 0.950 0.188 4.1 4.9 350
QUEEN-s [girish2025queen] 31.14 0.954 0.173 0.45 3.8 421
ComGS-s (ours) 31.49 0.955 0.171 0.028 28.3 98
Table 2: Quantitative comparisons on the long video sequence flame salmon from the N3DV dataset [li2022neural].
Method PSNR (dB) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Storage (MB) $\downarrow$ Training (sec) $\downarrow$ Rendering (FPS) $\uparrow$
E-NeRF [lin2022efficient] 23.48 0.89 0.260 0.692 13.8 5
4DGS [yang2023real] 28.89 0.952 0.197 2.23 31.2 90
TGH [xu2024representing] 29.44 0.945 0.214 0.075 6.3 550
ComGS-s (ours) 29.56 0.920 0.140 0.053 45.4 91
Figure 3: Quantitative comparison. We visualize our method and other online FVV methods on N3DV [li2022neural] and MeetRoom [li2022streaming] dataset.
Table 3: Ablation study on proposed components. Flame Steak and Flame Salmon scenes are from the N3DV dataset.
@@
[…]
@@ 4.1 Experimental Setup
We evaluate our method on two widely-used public benchmark datasets. (1) Neural 3D Video (N3DV) [li2022neural] consists of six indoor video sequences captured by 18 to 21 viewpoints. (2) Meet Room [li2022streaming] comprises four indoor scenes recorded with a 13 cameras multi-view system. In both of two datasets, we employ the first view for testing. Our method is implemented on an NVIDIA A100 GPU. We train 150 epochs for non-key frames reconstruction and 1000 epochs for key frames fine-tuning. We measure the visual quality of rendered images by average PSNR, required storage, rendering FPS and training time. More implement details are provided in the Appendix.
@@ 4.2 Quantitative Comparisons
We conduct quantitative comparisons on existing online methods including StreamRF [li2022streaming], TeTriRF [wu2024tetrirf], 3DGStream [sun20243dgstream] and QUEEN [girish2025queen], as well as the SOTA offline FVV approaches [li2024spacetime, wu20244d, song2023nerfplayer, attal2023hyperreel] on N3DV and Meetroom (Tab. 1). Our method is evaluated in two variants: ComGS-s (small) and ComGS-l (large), using key frame intervals of s=10s{=}10 and s=2s{=}2, respectively.
Tab. 1 shows that our ComGS achieves competitive results among existing online FVV methods on N3DV dataset. Notably, ComGS-s achieves a substantial reduction in storage by $159\times$ compared to 3DGStream and $14\times$ compared to QUEEN. This advantage enables real-time transmission in limited bandwidth and enhances the overall user viewing experience. On MeetRoom dataset, our method outperforms 3DGStream, obtaining +0.7dB PSNR and 146 $\times$ smaller size. Our advantages are mainly due to two factor: 1) using keypoint as a shared representation requires transmitting only a small number of keypoint attributes; and 2) the error-aware correction module effectively rectifies regions with scene inaccuracies using minimal additional parameters. In the Appendix, the quantitative results are provided for each scene to offer a more detailed comparison.
We further evaluate the effectiveness of ComGS on handling long videos. We compare our method with the TGH [xu2024representing] (which is the most recently proposed for handling long video sequences) on the Flame Salmon sequence (1200 frames) from the N3DV dataset [li2022neural]. Tab. 2 shows that our method achieves competitive results on rendering quality and required storage.
@@ 4.4 Ablation Study
To validate the effectiveness of our proposed methods, we ablate three components of ComGS framework in Tab. 3.
In the Experiment 1, we ablate the motion-sensitive keypoint selection and instead select keypoints randomly. Since the random selection is not guided by motion regions, it may result in ineffective modeling of static areas and inadequate representation on motion regions (Fig. 5 (b)), which leads to a slight degradation in PSNR. Experiment 2 removes the adaptive motion-driven mechanism and models scene motion only using the selected keypoints, without incorporating any neighboring points. The resulting drop in reconstruction quality demonstrates that effective motion modeling relies not only on accurately keypoint selection, but also on the selection of their neighboring points. In the Experiment 3, we reconstruct FVV only relying on the first frame reconstruction and key frame correction, without modeling non-key frames by keypoint reconstruction, which results in a significant performance drop. We emphasize that although the parameters of keypoints are few, the keypoint-based modeling plays a crucial role in FVV reconstruction. Experiment 4 ablates the error-aware correction in key frame reconstruction. The performance degradation demonstrates that the error-aware correction in key frames would solve the error accumulation across frames.
@@
To further investigate the role of keypoint-driven motion representation, we visualize the selection and driven process in Fig. 4. The top row shows that keypoints are predominantly selected in motion regions, such as the human body and moving objects. The bottom row highlights the adaptively controlled areas for neighboring points, which similarly focus on regions with significant motion (e.g., the person and the dog). Fig. 5 visualizes Gaussians updated region using farthest keypoint selection [wan2024superpoint], random keypoint selection and our method, respectively, which demonstrates that our method accurately captures motion-intensive areas. These results indicate that ComGS can effectively leverage the locality and consistency of scene motion.
We also evaluate a KNN-based method [peterson2009k] for selecting neighboring points around keypoints (Tab. LABEL:Tab:_ablation_knn). This approach shows inferior performance, as it does not distinguish between static and motion regions, leading to redundant modeling and poor adaptation to varying motion scales.